Acct 2302
19. The following materials standards have been established for a particular product: Standard quantity per unit of output Standard pric. 28 grams $12.50 per gram The following data pertain to operations concerning the product for the last month Actual materials purchased6,200 grams Actual cost of materials purchased..$81,530 Actual materials used in production... 5,700 grams Actual outpu800 units What is the materials price variance for the month? A. $6,250 U B. $4,030 U C. $8,679 U D. $6,575 U 20. Dube Corporation is considering the following three investment projects Project D Project E Project Investment required Present value of cash inflows $11,000 $41,000 $86,000 $11,330 $46,330 $95,4 Rank the projects according to the profitability index, from most profitable to least profitable. A. F.ED B. D.F.E C.F.D.E D. E,F,D use in four years. How much should Mary invest now in order to have the $20,000 available in four years if she can invest money at 16%: A. $13,990 B. $2,760 C. $11,040 D. $16,800 22. An opportunity cost mains the same regardless of the altemative chosen. B. Requires a current outlay of cash. C. Results from past managerial decisions. D. Is the potential benefit lost by choosing a specific alternative course of action among two or more E. Is irrelevant in decision making because it occurred in the past.
Answer #1

QUESTION - 19

Material Price Variance = Actual Quantity x [Actual rate – Standard rate]

= AQ x [AR – SR]

= 6,200 x [$13.15 – 12.50]

= 4,030 U [Unfavorable]

*AR = $81,530 / 6,200 grams = $13.15

QUESTION - 20

Profitability Index = Present Value of Annual Inflows / Investment Required

Project D = $11,130 / 11,000 = 1.03 - RANK 3

Project E = $46,330 / 41,000 = 1.13 – RANK 1

Project F = $95,460 / 86,000 = 1.11 - RANK 2

Therefore, The Correct Answer option is “D. E,F,D”

QUESTION – 21

Present Value = Future Value / (1+r)n

Present Value = $20,000 / (1+0.16) 4

Present Value = $20,000 / 1.81063936

Present Value = $11,040

Hence, The Answer is “C. $11,040”

QUESTION – 22

An Opportunity cost “is the potential benefit lost by choosing a specific alternative course of action among two or more”